2) C (3, -3), D (0, 4) Find the midpoints of the ff. Points 3) E (3, 2) , F (-1 , -2) 4) G (-3, 0), H (2, -1 ) 5) I (3, 1), J (-2, 0 )


1
Expert's answer
2022-02-22T11:13:11-0500

2)


"x_M=\dfrac{x_C+x_D}{2}=\dfrac{3+0}{2}=3/2"

"y_M=\dfrac{y_C+y_D}{2}=\dfrac{-3+4}{2}=1/2"

"M(3\/2, 1\/2)"


3)


"x_K=\dfrac{x_E+x_F}{2}=\dfrac{3-1}{2}=1"

"y_K=\dfrac{y_E+y_F}{2}=\dfrac{2-2}{2}=0"

"K(1, 0)"


4)


"x_N=\dfrac{x_G+x_H}{2}=\dfrac{-3+2}{2}=-1/2"

"y_N=\dfrac{y_G+y_H}{2}=\dfrac{0-1}{2}=-1/2"

"N(-1\/2, -1\/2)"


5)


"x_P=\dfrac{x_I+x_J}{2}=\dfrac{3-2}{2}=1/2"

"y_P=\dfrac{y_I+y_J}{2}=\dfrac{1+0}{2}=1/2"

"P(1\/2, 1\/2)"